# Rydberg Atomic Quantum Receivers for Classical Wireless Communication and Sensing

### v1

@@ IV Rydberg Atomic Quantum SISO and MIMO Schemes
In this section, we introduce the RAQRs into classical wireless systems by detailing the transceiver structure of the RAQ-SISO system, and presenting a potential RAQ-MIMO scheme, which evolved from the schemes portrayed from a physics perspective in [2, 7].
IV-A RAQ-SISO Architecture
Fig. 4 shows each stage of the holistic RAQ-SISO system from a communication perspective. At the transmit side, a classic wireless transmitter converts the baseband signal to a passband RF signal by the following two blocks:
  1. TXB1
    Interpolation: Following Nyquist sampling, the transmitted baseband signal (assuming a bandwidth of $W$) is transformed to its continuous-time counterpart using the sinc function.
  2. TXB2
    Up-conversion: The continuous-time baseband signal of zero center frequency is then upshifted to a passband RF signal of carrier frequency $f_{c}$.
As indicated by the downwards dotted arrows of Fig. 4, the details of the two transmit blocks are seen in the dotted boxes, respectively, where the transmitter relies on an in-phase and quadrature-phase (IQ) modulation scheme, and mainly operates in the digital and RF domains.
The passband RF signal propagates through the wireless channel characterized by a linear system response $h(\tau,t)$, and it is detected by the RAQR based wireless receiver including:
  1. RXB1
@@
    Rydberg Atomic sensing: It realizes RF reception using the superheterodyne RAQR that outputs an optical signal, which can be approximated by an amplified cosine wave of frequency $f_{c}-f_{l}$ [6].
  2. RXB2
@@
    Photodetection: The optical signal is then detected by a photodetector, which can be a direct incoherent optical detection (DIOD) scheme or a balanced coherent optical detection (BCOD) scheme. Both DIOD and BCOD generate an output that can be approximated by a cosine wave of frequency $f_{c}-f_{l}$.
  3. RXB3
@@
    Down-conversion: To obtain a continuous-time baseband signal of the photodetector’s output, down-conversion is realized by a lock-in amplifier, which requires IQ mixers having a mixing frequency of $f_{c}-f_{l}$ and lowpass filters having a bandwidth of $W$.
  4. RXB4
@@
    Sampling: The continuous-time baseband signal is then sampled and quantized by analog-to-digital converters (ADCs) to acquire the baseband samples. The sampling rate obeys Nyquist’s sampling theorem.
  5. RXB5
@@
    Baseband recovery: The baseband samples embody both the intensity and phase of the desired RF signal, which are ultilized for recovering the transmitted baseband signal using specific signal processing algorithms.
The receiver blocks are detailed in their dotted boxes, respectively, where the RAQR based wireless receiver operates in the optical, RF and digital domains.
This architecture is capable of integrating RAQRs into a holistic classical communication system, and paving the way for constructing the corresponding mathematical models and signal processing algorithms.
Figure 5: RAQ-MIMO receiver.
IV-B RAQ-MIMO Architecture
To extend the RAQ-SISO of Fig. 4 to a MIMO architecture, the most straightforward way is to use multiple vapor cells to form an array, where each vapor cell is penetrated using a probe-coupling beam pair to form the Rydberg states for RF reception, as seen in Fig. 4. The optical outputs of all vapor cells are forwarded to the corresponding photodetector, lock-in amplifier and ADC. The outputs of all branches are then appropriately combined to recover the desired RF signal, as seen in Fig. 5(a).
In addition to the above MIMO extension, we introduce a more compact one, as presented in Fig. 5(b). This architecture is implemented using a single vapor cell penetrated by a laser beam array. The optical array signals experience diversity gains owing to their different locations, and are detected by a photodetector array. The outputs are then delivered to multiple lock-in amplifiers and ADCs for baseband recovery. This architecture of the RAQ-MIMO receivers is initially verified in [7] to realize angle-of-arrival detection. This scheme is expected to play an important role in future quantum-aided wireless MIMO systems given its compact form factor.
@@ V Open Problems and Future Directions
Even though significant advances have been achieved in the evolution of RAQRs, the answers to several open problems remain to be further explored. We thus list a number of future directions of significant interest as follows. Advanced designs of RAQR have the potential of approaching or even exceeding the SQL, thereby outperforming the Shannon-limit of the current RF receivers.
Theoretical Modeling of RAQR-Aided Wireless Systems: At the time of writing, the integration of RAQRs into communication and sensing systems is in its infancy. The theoretical modelling is mainly dedicated to the input-output relationship of the vapor cell. However, a holistic baseband equivalent mathematical model, especially the baseband equivalent model, has to be developed for the RF signals of various systems (e.g., SISO, MIMO).
Signal Processing of RAQR-Aided Wireless Systems: Furthermore, computationally efficient signal processing algorithms must be conceived for communication-/sensing-oriented functionalities, including signal detection, channel estimation, and spatial displacement/direction detection etc.
Experimental Implementations and Validations: Previous efforts have shown the potential of RAQRs, but substantive future efforts are required for experimentally investigating its practical capability. Furthermore, the attainable system-level benefits must be quantified.
Cost, Size, Weight, and Power Reduction: To commercialize RAQRs, mature cost-saving, miniaturization and power-conserving technologies of the laser generators and their accessories are required to realize such receivers.
@@ VI Conclusions
RAQRs and their integration into wireless communication and sensing were considered. We briefly introduced the fundamentals of Rydberg atoms and discussed the RAQRs’ capabilities as well as distinctive features, followed by a contemporary survey of the state-of-the-art studies. To support the integration of RAQRs with classical wireless systems, we advocated RAQ-SISO and RAQ-MIMO schemes. Finally, we concluded by proposing promising research directions.

### v2

@@ IV RAQ-SISO and RAQ-MIMO Schemes
In this section, we introduce the RAQRs into classical wireless systems by detailing the receiver structure of RAQ-SISO systems, and presenting potential RAQ-MIMO schemes in both centralized and distributed applications.
IV-A Reception Scheme of RAQ-SISO Systems
Fig. 4 shows each stage of the holistic RAQ-SISO system and its corresponding signal model from an equivalent baseband communication perspective. The received input signal of the RAQR may be a communication signal from a classical wireless transmitter or a sensing signal reflected from some moving target. The received signal is assumed to have a carrier frequency of $f_{c}$, which is processed by the following blocks.
  1. B1
@@ IV-A Reception Scheme of RAQ-SISO Systems
    Rydberg atomic quantum sensing: It realizes RF reception using superheterodyne RAQRs. Its optical output is modulated and can be approximated by an amplified cosine wave of frequency $f_{c}-f_{l}$ [8, 18].
  2. B2
@@
    Photodetection: The optical signal is detected by a PD, which can be realized by direct incoherent optical detection (DIOD) or balanced coherent optical detection (BCOD) schemes [18]. Both schemes generate an output that is approximated by a cosine wave of frequency $f_{c}-f_{l}$.
  3. B3
@@
    Down-conversion: To obtain a continuous-time baseband signal of the PD’s output, down-conversion is realized by a homodyne receiver (HR), which requires IQ mixers having a mixing frequency of $f_{c}-f_{l}$ and lowpass filters having a bandwidth of $W$ (nearly the same as the signal bandwidth). We emphasize that this functionality can be implemented in either RF or digital domains.
  4. B4
@@
    Signal sampling: The continuous-time baseband signal is sampled by analog-to-digital converters (ADCs) and then the complex baseband samples are obtained. The sampling rate obeys Nyquist’s sampling theorem.
  5. B5
@@
    Baseband signal processing: The baseband samples embody both the intensity and phase of the desired RF signal, which are utilized for recovering the transmitted baseband signal or for estimating the information of the target using specific signal processing algorithms.
The equivalent baseband signal model of the holistic wireless system constructed in [18] is briefly summarized at the bottom of Fig. 4, where the parameters are explained in the box at the bottom. It is seen that the RAQR imposes a gain $\varrho$ and a phase shift $\Phi$ on the received RF signal. These parameters are determined by both the quantum response and the photodetection scheme selected. The accuracy of this signal model has been verified in [18]. In our simulations of this article, we follow the same parameter configurations of the unoptimized RAQR and of the conventional RF receiver as in [18] and showcase the received SNR and bit-error-rate (BER) versus the transmit power under Rayleigh fading channels. The transmission distance is 200 meters having a path loss factor of 3.8. The results reveal that the RAQ-SISO system significantly outperforms the conventional SISO systems under the same transmit power, as seen in the simulation results of Fig. 4.
Figure 5: (a) Centralized RAQ-MIMO, (b) distributed RAQ-MIMO, and (c) simulation results of the centralized RAQ-MIMO.
IV-B Centralized and Distributed RAQ-MIMO Schemes
IV-B1 Centralized RAQ-MIMO
The most straightforward way to form a co-located array is using multiple vapor cells, where each vapor cell is penetrated using a probe-coupling beam pair to form the receive region, as seen at the top of Fig. 5(a). Furthermore, we introduce a more compact one shown at the bottom of this figure, which is implemented by penetrating a single vapor cell using a laser beam array. The former scheme has the advantage of constructing large receiver apertures, while the latter scheme facilitates compact receiver apertures. For both schemes, their parallel optical outputs are forwarded to the corresponding photodetector array (PDA), HRs, and ADCs. Therefore, all samples can be jointly modelled to obey a corresponding signal model. These samples are then appropriately combined to realize the desired communication/sensing tasks by using specific algorithms, as seen in the ‘Signal model’ and ‘Performance analysis & signal processing’ boxes of Fig. 5(a).
For both of the above schemes, the optical array signals experience diversity gains owing to their different locations. Compared to the OLA scheme [9], these centralized RAQ-MIMO schemes combine all branch signals in the digital domain instead of directly combining all the output optical signals. This retains higher degrees-of-freedom for the information of signal and facilitates better performance. Furthermore, we present simulation results of a single vapor cell having a uniform linear array of 5 laser beams (5 receive regions), where each region is identically configured as that of the RAQ-SISO. The achievable rate of wireless communication and the Cramér–Rao bound (CRB) of DOA estimation are showcased for a single user/target at a distance of 200 meters. As observed from the simulation results of Fig. 5(a), the achievable rate of RAQ-MIMO is $\sim 7.2$ bits/s/Hz higher than that of the conventional MIMO systems under Rayleigh fading channels. Additionally, the CRB achieved by the RAR-MIMO improves that achieved by the conventional MIMO by as much as two orders of magnitude.
IV-B2 Distributed RAQ-MIMO
To further enhance the spatial diversity and improve the coverage, we present the distributed RAQ-MIMO scheme, where the system consists of multiple vapor cells and a single data center. Specifically, the vapor cells are spatially distributed across a long distance range, so that each vapor cell is capable of receiving RF signals having larger spatial diversity or covering more users/targets. The data center consists of laser generators, the PDA/multiple PDs (responsible for receiving long-distant optical signal), the HRs, ADCs, and the distributed signal processing unit.
The data center and vapor cells are connected in two ways, namely the fiber-coupled and optically-linked schemes of Fig. 5(b). These schemes are based on the recent advances of fiber-coupled vapor cells [2] (references [117], [119] therein) and of passive RAQRs at long distances [20]. The two schemes have their own advantages and shortcomings. Specifically, the fiber-coupled scheme appear to be more robust to environment uncertainty than the opticall-linked scheme due to the higher reliability of fibers, while the latter exhibits higher mobility and is easier to be deployed.
@@ V Open Problems and Future Directions
Advanced designs of RAQRs have the potential of approaching or even exceeding the SQL, thereby outperforming the Shannon-limit of the current RF receivers. Even though significant advances have been achieved in the evolution of RAQRs, the answers to several open problems remain to be further explored. Therefore, we list a number of future directions of significant interest as follows.
Improvements of Receiver Metrics and Commercialization of RAQRs: Albeit the basic capabilities of RAQRs have been explored and verified. Future studies are required to be continued for pushing forward the receiver performance, such as the sensitivity, bandwidth, and robustness, etc. Furthermore, to commercialize RAQRs, mature cost-saving, miniaturization, and power-conserving technologies are required.
Applications of RAQRs to Next-Generation Wireless Communication and Sensing: To fully harness the benefits of RAQRs, their integration into wireless communication and sensing is high of significance. At the time of writing, the integration is in its infancy. More particularly, we expect three main directions to be explored.
  1. D1
    Theoretical Modelling of RAQR-Aided Wireless Systems: The currently available theoretical modelling is mainly dedicated to the input-output relationship of the vapor cell. However, the holistic baseband equivalent models have to be developed for various wireless systems (e.g., wideband, MIMO).
  2. D2
    Signal Processing of RAQR-Aided Wireless Systems: With the aid of mathematical models constructed, computationally efficient signal processing algorithms must be conceived for communication-/sensing-oriented functionalities, including signal detection, channel estimation, and spatial displacement/direction detection etc.
  3. D3
    Interplay between RAQRs and Classical Transmitters: In RAQR-aided wireless systems, classical antenna-based transmitters are still needed. The interesting interplay between RAQRs and classical transmitters opens a new research direction on the design of novel transceivers.
  4. D4
    System-Level Experimental Implementations and Validations: Previous efforts have shown the great potential of RAQRs, but substantive future efforts are required for experimentally investigating their practical system-level capabilities and benefits.
@@ VI Conclusions
In this article, RAQRs and their integration into wireless communication and sensing have been considered. We have briefly introduced the fundamentals of Rydberg atoms and discussed the RAQRs’ capabilities as well as distinctive features, followed by a contemporary survey of the state-of-the-art studies. To support the integration of RAQRs with classical wireless systems, we have advocated the RAQ-SISO and RAQ-MIMO schemes. Finally, we have concluded by proposing promising research directions.